Two men can swim in still water at the rate of 5 kmph. One swims upstream a distance equal to the width of the river (half a km) and back again; the other swims to a point straight across the river and back again. Assuming that the velocity of the river is 3 kmph, find the difference in times taken by two men. (Given answer is 3.75 min.)


I am getting a time difference of 1.8 min.


Please help

Grade:8

1 Answers

Ankit Khokhar
12 years ago

Sorry but your answer is wrong!!!!!!!

3.75 minutes is the correct answer.

for the first person, speed in first path=2 kmph (as he is swimming backward)

and in second path speed=8 kmph  (as he is swimming forward)

so total time taken by him= (1/2) /2 + (1/2)/8 = 5/16 hours

now for second person he should swim in a direction so that when seen from ground his speed should be perpendicular to the flow of river so he should swim at an angle x, where sin x=3/5. so his speed will be 4 kmph in direction perpendicular to the flow of river.

So time taken by second person= (1/2)/4 + (1/2)/4          (as condition is same in both cases of second person)

so total time taken by second=1/4 hours.

so difference in time= (5/16 - 1/4) =1/16 hours = 3.75 minutes.

Hope you understand!!!!!!!!!!!!